What is the probability of two people in our class having the same birthday? Briefly tell whether you think it will be a high probability or low (or somewhere in between) and then try to calculate!

How would I calculate this or answer this question
Answer #1

Solution :

\mathrm{The\:chance\:that\:the\:first\:person\:has\:the\:same\:birthday\:is\:1/365,}

\mathrm{The\:chance\:that\:the\:second\:person\:has\:the\:same\:birthday\:is\:1/365,}

\mathrm{The\:probability\:that\:both\:people\:have\:this\:birthday,\:for\:that\:we\:have\:to\:multiply\:their\:separate\:probabilities}

=\left (\frac{365}{365} \right ) * \left (\frac{1}{365} \right )

=\frac{1}{365}

\mathrm{Hence,}

\mathrm{The\:chance\:of\:the\:first\:and\:second\:person\:sharing\:a\:birthday\:is\:still\:\left ( \frac{1}{365} \right ).}